OpenStudy (anonymous):

How do you solve 2(x-3)^2=98 algebraically? My answer key says the answer is 10, but I think it is wrong

OpenStudy (anonymous):

No, its right, Start by dividing both sides by 2, then you get (x-3)^2=49 Then square root both sides and you get (x-3)=7 Then add 3 to both sides and you get x=10

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

Anytime :)

OpenStudy (nikvist):

\[2(x-3)^2=98\] \[(x-3)^2=49=7^2=(-7)^2\] \[x-3=7\quad\vee\quad x-3=-7\] \[x=10\quad\vee\quad x=-4\]